. Compute the x- and y-components of the vectors.. ,A=8.00m, B=15.00m, C=12.00m and D=10.0
1
Expert's answer
2019-10-11T10:36:45-0400
"A_x=0\ m, A_y=-8\ m"

"B_x=15\sin{30}=7.5\ m, B_y=15\cos{30}=12.99\ m"

"C_x=-12\cos{25}=-10.88\ m, C_y=-12\sin{25}=-5.07\ m"

"D_x=-10\sin{53}=-7.99\ m, D_y=10\cos{53}=6.02\ m"